OpenStudy (anonymous):

John can do a piece of work in 8 days. If Paul can do it in 10 days, in how many days can both do it together? Please explain step by step (and explain) :D

OpenStudy (anonymous):

john does 1/8th a pieace in a day, paul does 1/10 a piece in a day. the equation to solve is x((1/8)+(1/10))=1 where x is the number of days and it equals 1 because you're looking for a complete piece of work

OpenStudy (anonymous):

and the answer is....ba badum dum....(insert dramatic music).....

OpenStudy (anonymous):

x = 1/((1/8)+(1/10)) = 40/9

OpenStudy (anonymous):

is it 40/9 or 4 (fraction bar) 4/9

OpenStudy (anonymous):

40/9 = 4+(4/9) = 4.444

OpenStudy (anonymous):

Thanks :D